OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the third equation. 2x – 3y + z = –4 2x – 2y + z = –1 x – 2y + 3z = –6 Answer –y – 1 = –4 –5x + 4y = –3 y – 5z = 8 2y – 5z = 11 5x + 7y = 6 y – 4 = –1 –7y + 7z = 8 –6y + 7z = 11

OpenStudy (anonymous):

i need help x+y=-2, 2x-2y=-8 (-3 1) substitution method

OpenStudy (anonymous):

x+y=-2, 2x-2y=-8 (-3 1) substitution method